# Effects of dynamical friction on perturbations for evolving dark energy

### v1

@@ Abstract
We explore the impact of dynamical friction on cosmological scales and show its influence on the evolution of perturbations. In particular, considering smooth and clustering dark energy models, we describe the role played by friction by selecting two main hierarchical models, i.e., the first where the friction term is proportional to the Hubble rate, whereas the second where friction is induced by the dark energy pressure. The second approach generalises the first and translates the idea that pressure is a general relativistic effect, motivating why friction might arise once barotropic dark energy fluids are considered. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, using the formalism of the spherical collapse model. Whilst dynamical friction has very small effects and thus it cannot be excluded a priori, dissipative pressure friction leads to a substantial slow down in the evolution of perturbations. This can be particularly inferred within the halo mass function, for which we also employ corrections due to dark energy clustering. To this end, in order to discern detectable deviations from the standard cosmological model, we thus highlight where dissipation effects might play a significant role at large scales.
pacs:
98.80.Jk, 98.80.-k, 98.80.Es
I Introduction
@@
Ever since the discovery of cosmic acceleration Riess et al. (1998); Perlmutter et al. (1998), several probes undoubtedly showed that the universe expansion rate is currently speeding up, through something unknown dubbed dark energy, or in the simplest case by vacuum energy Martin (2012). In order to disclose the dark energy nature, several theoretical models have been proposed throughout the last two decades Copeland et al. (2006).
@@
Besides the existence of low- and high-redshift tests Aviles et al. (2012), alternatives can make use of morphological features of the observable universe, among which high-precision galaxy clustering and weak gravitational lensing measurements, to check how dark matter can cluster. The clustering process is intimately related to small perturbations induced by cosmic fluids in the universe and, in fact, any dark energy model might overcome successfully the aforementioned low- and high-redshift tests, together with morphological probes.
@@
In this respect, the concordance background paradigm, i.e., the best candidate to explain how the acceleration occurs, dubbed the $\Lambda$CDM model Bull et al. (2016), where $\Lambda$ represents the cosmological constant and CDM the cold dark matter component, unequivocally appears to be the best description of the Cosmic Microwave Background (CMB) and Large Scale Structure (LSS) surveys. Phrasing it differently, the concordance approach depends on the fewest number of free parameters possible Perivolaropoulos and Skara (2022), suggesting that any dark energy model is statistically disfavoured with respect to it [^1] For the sake of clarity, the $\Lambda$CDM model requires six parameters, albeit at late-times, one single parameter, namely the mass density, approximates well the overall dynamics. Thus, to falsify the model, a compelling dark energy scenario might be described by one parameter only. The only class of dark energy models prompting one parameter only, being different from $\Lambda$CDM is the dark fluid, see e.g. Arbey (2006); Aviles and Cervantes-Cota (2011); Luongo and Muccino (2018); D’Agostino et al. (2022). Model-independent tests seem to certify the goodness of the standard background model at both low and intermediate levels Luongo and Muccino (2020). Shafieloo et al. (2018).
@@
So, apparently, there is no reason to go further as the $\Lambda$CDM model describes the universe so well at large scales. However, even though the cosmological constant $\Lambda$, within the context of general relativity, describes perfectly well the cosmic speed up, a new dark sector physics seems to be required in order to heal current tensions Di Valentino et al. (2021); Douspis et al. (2018) and the associated caveats related to the $\Lambda$CDM model itself.
@@
Consequently, the complete understanding of nonlinear structure formation becomes central to model the power spectrum and then to infer deviations from the concordance paradigm. In this respect, rich clusters can be used to constrain cosmological models as they represent the largest virialized objects in the universe [^2] Hence, they can be put in relation with the linear matter power spectrum, whereas their temperature can be used to infer the cluster mass..
@@
Several alternatives to the standard concordance model have been investigated Copeland et al. (2006); Capozziello et al. (2019a), with models spanning from extending general relativity, up to barotropic dynamical dark energy, passing through interacting dark energy models [^3] Most of these approaches assume some form of energy transfer that affects both background and perturbations. These models usually fail to fit CMB data or are severely constrained..
@@
Motivated by the above points, we here investigate the effects due to dynamical friction at large scales. In this respect, we consider both smooth and clustering dark energy models, limiting the analysis, for the sake of simplicity, to models with a linearly evolving equation of state Chevallier and Polarski (2001); Linder (2003). We choose this class of models to guarantee observations to hold, as dark energy scenarios seem to favour a weakly-evolving dark energy term, instead of more complicated functions [^4] For these reasons, we exclude extended theories of gravity and we focus on general relativity only. Cosmic tests seem to agree with this assumption Calzà et al. (2020); Muccino et al. (2021). Capozziello et al. (2014); Gruber and Luongo (2014); Capozziello et al. (2015); Dunsby and Luongo (2016); Aviles et al. (2014); Capozziello et al. (2020); Aviles et al. (2013a, b); Luongo (2011); de la Cruz-Dombriz et al. (2016); Capozziello et al. (2019b); Capozziello et al. (2018); Luongo (2013); Aviles et al. (2017). In particular, to describe the role played by friction, we single out two main models. In the first scenario, friction is proportional to the Hubble rate, whereas the second shows a more complicated correction that is mostly proportional to the dark energy pressure. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, adopting the basic demands of spherical collapse theory. The evolution of perturbations is studied and the initial conditions properly inferred. We display our numerical bounds, showing that the main discrepancy occurs for the first hierarchy, at both the levels of linear and nonlinear perturbations. Consequences on the large scale structures induced by the presence of friction are reported in detail, showing the main differences to standard frictionless cases.
@@
Our numerical findings indicate that there is no reason to exclude a priori the presence of friction in cosmological realms. Even though the standard cosmological model does not take into account dissipation effects, we predict suitable bounds within the presence of dissipation can in principle be introduced. Since this property would significantly affect the overall clustering, we propose the viable ranges to detect it. Our numerical results suggest that there is no precise justification to exclude the presence of friction in cosmological realms a priori. Even though the standard cosmological model neglects dissipation effects, we propose potential limits of friction within clustering. To do so, given the considerable impact of this recipe in structure formation, we propose feasible ranges for its detection, with the aim to detect dissipation effects in cosmology.
@@
The paper is organised as follows. In Sect. II we present a detailed derivation of the equations of motion for the collapsing fluid and in Sect. III we discuss the models used in this work. In Sect. IV we discuss the evolution of linear perturbations while the study of nonlinear perturbations is presented in Sect. V. In Sect. VI we show the effects of dynamical friction on the halo mass function. We conclude our work in Sect. VII.
II Extended spherical collapse model with friction
@@
In this section we derive the expressions needed to study the evolution of linear and nonlinear perturbations of dark matter in a dark energy model where the dark energy component is smooth and only affects the background evolution. We will then extend the formalism to include perturbations also for dark energy.
@@
The spherical collapse model assumes that perturbations are spherically symmetric and isolated Tomita (1969); Gunn and Gott (1972); Bernardeau (1994); Lee and Ng (2010). A further assumption, which allows to greatly simplify the equations, is that perturbations are described by a top-hat profile, i.e., the profile does not change in space but it is constant over the size of the perturbation. The value of the density can, therefore, only change in amplitude, but the shape of the profile stays the same. This is quite a crude approximation, but it has been proven to be very effective in describing the nonlinear evolution of perturbations, showing a good agreement between the theoretical and the numerical mass function in $N$-body simulations for both dark energy Pace et al. (2010) and modified gravity models Pace et al. (2014a).
@@
This basic picture has been improved in several ways, mainly relaxing the assumption of spherical symmetry and isolated perturbations. Ohta et al. (2003, 2004) assumed triaxial (ellipsoidal) perturbations and took into account tidal fields; Del Popolo et al. (2013a, b, c); Pace et al. (2014b) considered the effects of shear and rotation (we will define them in a precise way later) assuming an expression derived by modelling the angular momentum of galaxies and clusters while Reischke et al. (2016); Pace et al. (2017a); Reischke et al. (2018) proposed an approach based on the Zel’dovich approximation. Finally, Pace et al. (2019); Pace and Schimd (2022) used deviations from sphericity to study the virialization of cosmic structures and Abramo et al. (2007, 2008); Chang et al. (2018) considered perturbations also in the dark energy component.
@@
Here, we build upon all these works by further extending the formalism to include the effects of dynamical friction in $\Lambda$CDM and dark energy models. Hence, we will derive the equations of motion for matter perturbations taking into account shear, rotation and dynamical friction.
@@
Dynamical friction manifests itself during the evolution of perturbations and affects their evolution slowing down their collapse. This means that cosmic structures require more time to form. As in Newtonian dynamics, friction affects the equations of motion and, being proportional to velocity, it modifies the Euler equation, which now reads (Carrilho et al., 2022)
@@
In the previous equation, $H$ is the Hubble function, $\mathbf{u}$ represents the comoving peculiar velocity defined as $\mathbf{v}=aH\mathbf{x}+a\mathbf{u}$, with $a$ the scale factor, $\mathbf{x}$ the comoving position and $\mathbf{v}$ the fluid velocity. The perturbed potential is denoted by $\phi$. Finally, $\eta=\eta(t)$ is an arbitrary function of time denoting the evolution of dynamical friction. As it is evident from Eq. (1), dynamical friction has dimensions of an inverse of time.
@@
with $\Omega_{\rm m}(a)$ the matter density parameter, $\delta=\delta\rho/\bar{\rho}$ the fluid density perturbation and the prime represents the derivative with respect to $\ln{a}$.
@@
with $\theta=\nabla\cdot\mathbf{u}$ the divergence of the peculiar velocity and $\delta_{ij}$ the Kronecker’s delta symbol.
If we linearise Eq. (II), we obtain the modified equation describing the evolution of the growth factor:
@@
Both the linear and nonlinear expressions assume the same form of those obtained for dark scattering models Baldi and Simpson (2017); Carrilho et al. (2021, 2022). In these models there is an interaction between the cold dark matter and the dark energy component which gives rise to an exchange of energy between the two fluids. The equations of motion lead to an additional friction term which modifies the coefficient of $\delta^{\prime}$ and depends on the cross section and mass of the dark matter particles. Here we want to move a step further and study the physical consequences of different implementations of the dynamical friction on the nonlinear evolution of perturbations, in particular on the mass function.
The formalism is easily extended to include the effects of clustering dark energy. To focus on this, let us now consider the following perturbation equations for the dark energy component (Abramo et al., 2007, 2008; Abramo et al., 2009a, b; Pace et al., 2017b):
$\delta_{\mathrm{de}}^{\prime}+3(s_{\rm eff}-w_{\mathrm{de}})\delta_{\mathrm{de}}+$
@@
$\tilde{\theta}_{\mathrm{de}}^{\prime}+\left(2+\frac{H^{\prime}}{H}\right)\tilde{\theta}_{\mathrm{de}}+\frac{\tilde{\theta}_{\mathrm{de}}^{2}}{3}+$
@@
where the prime implies again the derivative with respect to $\ln{a}$. Clearly, $\tilde{\theta}_{\mathrm{de}}=\theta_{\mathrm{de}}/H$ and $w_{\mathrm{de}}$ is the dark energy equation of state at the background with the corresponding sound speed defined by $c_{\rm eff}^{2}=c^{2}s_{\rm eff}=\delta P_{\mathrm{de}}/\delta\rho_{\mathrm{de}}$, where the speed of light, $c$, appears for dimensional reasons. This quantity is zero for the standard $\Lambda$CDM model and for classes of dark fluids, while it is fixed to $s_{\rm eff}=1$ for Quintessence. The different values that this can acquire have huge consequences in cosmology: a small sound speed implies structures to form at several scales, while high sound speeds correspond to stiff matter, suggesting perturbations to be strongly suppressed, i.e., involving dark energy that affects the background only. Consequently, it is licit to presume $s_{\rm eff}\ll 1$, guaranteeing that dark energy perturbations become relatively important in analogy to matter perturbations, albeit always subdominant (Batista and Pace, 2013). Note that, in general, $\theta_{\rm de}\neq\theta_{\rm m}$, as friction only affects (cold dark) matter.
II.1 A comment on dynamical friction
At this point, a deeper discussion of the results obtained is necessary. It is well known that an alternative way to discuss cosmological perturbation is based on the evolution of the physical radius of the perturbation $R$. In fact, many authors already considered the effect of dynamical friction from this point of view. The evolution of the (effective) radius of the perturbation $R$, including angular momentum $L$ and dynamical friction $\eta$, in a $\Lambda$CDM model is given by Lahav et al. (1991); Bartlett and Silk (1993); Peebles (1993); Antonuccio-Delogu and Colafrancesco (1994); Del Popolo and Gambera (1998, 1999) [^5] In these references, only single parts of the full expression are investigated or reported.

### v2

@@ Abstract
We explore the impact of dynamical friction on scales where the linear growth factor and the spherical collapse model can be applied and show its influence on the evolution of perturbations. In particular, considering smooth and clustering dark energy models, we describe the role played by friction by selecting two main hierarchical models, i.e., the first where the friction term is proportional to the Hubble rate, whereas the second where friction is induced by the dark energy pressure. The second approach generalises the first and translates the idea that pressure is a general relativistic effect, motivating why friction might arise once barotropic dark energy fluids are considered. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, using the formalism of the spherical collapse model. Whilst dynamical friction has very small effects and thus it cannot be excluded a priori, dissipative pressure friction leads to a substantial slow down in the evolution of perturbations. This can be inferred particularly from the halo mass function, for which we also employ corrections due to dark energy clustering. To this end, in order to discern detectable deviations from the standard cosmological model, we thus highlight where dissipation effects might play a significant role at large scales.
keywords:
Cosmology , dark energy , dynamical friction , spherical collapse
[…]
[1]organization=Dipartimento di Fisica, Università degli Studi di Torino,addressline=Via P. Giuria 1, postcode=I-10125, city=Torino, country=Italy
[…]
[11]organization=INFN-Sezione di Torino,addressline=Via P. Giuria 1, postcode=I-10125, city=Torino, country=Italy
[…]
[111]organization=INAF-Istituto Nazionale di Astrofisica, Osservatorio Astrofisico di Torino,addressline=strada Osservatorio 20, postcode=10025, city=Pino Torinese, country=Italy
[…]
[2]organization=Department of Mathematics and Physics, SUNY Polytechnic Institute, addressline=100 Seymour Rd, postcode=13502, city=Utica, state=New York, country=USA
[…]
[22]organization=Divisione di Fisica, Università degli studi di Camerino, addressline=Via Madonna delle Carceri 9B, postcode=62032, city=Camerino, country=Italy
[…]
[222]organization=INFN-Sezione di Perugia, addressline=Via Alessandro Pascoli 23c, postcode=06123, city=Perugia, country=Italy
[…]
[$2$]organization=Al-Farabi Kazakh National University, addressline=Al-Farabi Avenue 71, postcode=050040, city=Almaty, country=Kazakhstan
[…]
[3]organization=Dipartimento di Fisica e Astronomia, University of Catania, addressline=Viale Andrea Doria 6, postcode=95125, city=Catania, country=Italy
[…]
[33]organization=INFN-Sezione di Catania, addressline=Via Santa Sofia 64, postcode=95123, city=Catania, country=Italy
[…]
[333]organization=Institute of Astronomy, Russian Academy of Sciences, postcode=119017, addressline=Pyatnitskaya str., 48, city=Moscow, country=Russia
1 Introduction
@@
Ever since the discovery of cosmic acceleration [1, 2], several probes undoubtedly showed that the universe expansion rate is currently speeding up, through something unknown called dark energy, or in the simplest case by vacuum energy [3]. In order to disclose the dark energy nature, several theoretical models have been proposed throughout the last two decades [4].
@@
In addition to the existence of low- and high-redshift tests [5], alternatives can make use of morphological features of the observable universe, among which high-precision galaxy clustering and weak gravitational lensing measurements, to check how dark matter can cluster. The clustering process is intimately related to small perturbations induced by cosmic fluids in the universe, and, in fact, any dark energy model might successfully overcome the aforementioned low- and high-redshift tests, together with morphological probes.
@@
In this respect, concordance background paradigm, i.e., the best candidate to explain how the acceleration occurs, dubbed the $\Lambda$CDM model [6], where $\Lambda$ represents the cosmological constant and CDM the cold dark matter component, unequivocally appears to be the best description of the Cosmic Microwave Background (CMB) and Large Scale Structure (LSS) surveys. Phrasing it differently, the concordance approach depends on the fewest number of free parameters possible [7], suggesting that any dark energy model is statistically disfavoured with respect to it [8] [^1] For the sake of clarity, the $\Lambda$CDM model requires six parameters, albeit at late-times, one single parameter, namely the mass density, approximates well the overall dynamics. Thus, to falsify the model, a compelling dark energy scenario might be described by one parameter only. The only class of dark energy models that prompts one parameter only, being different from $\Lambda$CDM is dark fluid, see e.g. [9, 10, 11, 12]. Model-independent tests seem to certify the goodness of the standard background model at both low and intermediate levels [13]..
@@
So, apparently, there is no reason to go further, as the $\Lambda$CDM model describes the universe so well at large scales. However, even though the cosmological constant $\Lambda$, within the context of general relativity, describes well the cosmic speed up, new dark sector physics seems to be required in order to heal current tensions [14, 15] and the associated caveats related to the $\Lambda$CDM model itself.
@@
Consequently, a complete understanding of nonlinear structure formation becomes central to model the power spectrum and then to infer deviations from the concordance paradigm. In this respect, rich clusters can be used to constrain cosmological models, as they represent the largest virialized objects in the universe [^2] Hence, they can be put in relation with the linear matter power spectrum, whereas their temperature can be used to infer the cluster mass..
@@
Several alternatives to the standard concordance model have been investigated [4, 16], with models spanning from extending general relativity, up to barotropic dynamical dark energy, passing through interacting dark energy models [17, 18] [^3] Most of these approaches assume some form of energy transfer that affects both background and perturbations. These models often fail to fit the CMB data or are severely constrained..
@@
Motivated by the above points, we investigate here the effects resulting from dynamical friction at large scales. In this respect, we consider both smooth and clustering dark energy models, limiting the analysis, for the sake of simplicity, to models with a linearly evolving equation of state [19, 20]. We choose this class of models to ensure that the observations hold, as dark energy scenarios seem to favor a weakly evolving dark energy term, instead of more complicated functions [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] [^4] For these reasons, we exclude extended theories of gravity and we focus on general relativity only. Cosmic tests seem to agree with this assumption [35, 36].. In particular, to describe the role that friction plays, we have singled out two main models. In the first scenario, friction is proportional to the Hubble rate, whereas the second shows a more complicated correction that is mostly proportional to the dark energy pressure. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, adopting the basic demands of spherical collapse theory. The evolution of perturbations is studied and the initial conditions properly inferred. We display our numerical bounds, showing that the main discrepancy occurs for the first hierarchy, at both the levels of linear and nonlinear perturbations. The consequences on the large-scale structures induced by the presence of friction are reported in detail, showing the main differences to standard frictionless cases.
@@
Our numerical findings indicate that there is no reason to exclude a priori the presence of friction in the cosmological realms. Even though the standard cosmological model does not take into account dissipation effects, we predict suitable bounds within the presence of dissipation, which can in principle be introduced. Since this property would significantly affect the overall clustering, we propose viable ranges to detect it.
@@
The paper is organized as follows. In Sect. 2 we present a detailed derivation of the equations of motion for the collapsing fluid and in Sect. 3 we discuss the models used in this work. In Sect. 4 we discuss the evolution of linear perturbations, while the study of nonlinear perturbations is presented in Sect. 5. In Sect. 6 we show the effects of dynamical friction on the halo mass function. We conclude our work in Sect. 7.
2 Extended spherical collapse model with friction
@@
In this section, we derive the expressions needed to study the evolution of linear and nonlinear perturbations of dark matter in a dark energy model, where the dark energy component is smooth and affects only the background evolution. We will then extend the formalism to include perturbations also for dark energy.
@@
The spherical collapse model assumes that the perturbations are spherically symmetric and isolated [37, 38, 39, 40]. A further assumption, which allows to greatly simplify the equations, is that perturbations are described by a top-hat profile, i.e., the profile does not change in space but it is constant over the size of the perturbation. The value of the density can therefore only change in amplitude, but the shape of the profile stays the same. This is a crude approximation, but it has been proven to be very effective in describing the nonlinear evolution of perturbations, showing a good agreement between the theoretical and numerical mass functions in $N$-body simulations for both dark energy [41] and modified gravity models [42].
@@
This basic picture has been improved in several ways, mainly relaxing the assumption of spherical symmetry and isolated perturbations. [43, 44] assumed triaxial (ellipsoidal) perturbations and took into account tidal fields; [45, 46, 47, 48] considered the effects of shear and rotation (we will define them in a precise way later) assuming an expression derived by modeling the angular momentum of galaxies and clusters, while [49, 50, 51] proposed an approach based on the Zel’dovich approximation. Finally, [52, 53] used deviations from sphericity to study the virialization of cosmic structures and [54, 55, 56] considered perturbations also in the dark energy component.
@@
Here, we build upon all these works by further extending the formalism to include the effects of dynamical friction in $\Lambda$CDM and dark energy models. Hence, we will derive the equations of motion for matter perturbations taking into account shear, rotation, and dynamical friction.
@@
Dynamical friction manifests itself during the evolution of perturbations and affects their evolution, slowing their collapse. This means that cosmic structures require more time to form. As in Newtonian dynamics, friction affects the equations of motion and, being proportional to velocity, it modifies the Euler equation, which now reads
@@
In the previous equation, $H$ is the Hubble function, $\mathbf{u}$ represents the comoving peculiar velocity defined as $\mathbf{v}=aH\mathbf{x}+a\mathbf{u}$, with $a$ the scale factor, $\mathbf{x}$ the comoving position and $\mathbf{v}$ the fluid velocity. The perturbed potential is denoted by $\phi$. Finally, $\eta=\eta(t)$ is an arbitrary function of time that denotes the evolution of dynamical friction. As it is evident from Eq. (1), dynamical friction has dimensions of an inverse of time.
We note that this formulation, which later we will show to agree with the results derived based on the radius of the perturbation (upon some modifications we will discuss in detail), is very similar to that of models of dark scattering [57], but with an important difference. Whilst for dark scattering models the function $\eta$ is fixed by the theory, for models affected by friction this is a free function, whose specific form depends on the specific model chosen. As we shall discuss later, we shall consider two different functional forms, which are derived from different physical assumptions. Hence, despite the striking similarity to dark scattering models, here the phenomenology is much richer.
@@
with $\Omega_{\rm m}(a)$ the matter density parameter, $\delta=\delta\rho/\bar{\rho}$ the fluid density perturbation, $\tilde{\theta}=\theta/H$ with $\theta=\nabla\cdot\mathbf{u}$ being the divergence of the peculiar velocity and the prime represents the derivative with respect to $\ln{a}$.
@@
with $\delta_{ij}$ the Kronecker’s delta symbol.
If we linearise Eq. (2), we obtain the modified equation describing the evolution of the growth factor:
@@
Both the linear and nonlinear expressions assume the same form of those obtained for dark scattering models [58, 59, 57]. In dark scattering models, there is an interaction between the cold dark matter and the dark energy component, which gives rise to an exchange of energy between the two fluids. The equations of motion lead to an additional friction term which modifies the coefficient of $\delta^{\prime}$ and depends on the cross section and mass of the dark matter particles. However, we remark that the two models are physically distinct, and we mention dark-scattering models for completeness and for the similarity of the equations to be solved.
The formalism is easily extended to include the effects of clustering dark energy. To focus on this, let us now consider the following perturbation equations for the dark energy component [54, 55, 60, 61, 62]:
@@
where the prime implies again the derivative with respect to $\ln{a}$. Clearly, $\tilde{\theta}_{\mathrm{de}}=\theta_{\mathrm{de}}/H$ and $w_{\mathrm{de}}$ is the dark energy equation of state in the background with the corresponding sound speed defined by $c_{\rm eff}^{2}=c^{2}s_{\rm eff}=\delta P_{\mathrm{de}}/\delta\rho_{\mathrm{de}}$, where the speed of light, $c$, appears for dimensional reasons. This quantity is zero for the standard $\Lambda$CDM model and for classes of dark fluids, while it is fixed to $s_{\rm eff}=1$ for quintessence. The different values that this can acquire have huge consequences in cosmology: a small sound speed implies structures to form at several scales, while high sound speeds correspond to stiff matter, suggesting perturbations to be strongly suppressed, i.e., involving dark energy that affects the background only. Consequently, it is licit to presume $s_{\rm eff}\ll 1$, guaranteeing that dark energy perturbations become relatively important in analogy to matter perturbations, albeit always subdominant [63]. Note that, in general, $\theta_{\rm de}\neq\theta_{\rm m}$, as friction only affects (cold dark) matter.
2.1 A comment on dynamical friction
At this point, a deeper discussion of the results obtained is necessary. It is well known that an alternative way to discuss cosmological perturbations is based on the evolution of the physical radius of the perturbation $R$. In fact, many authors have already considered the effect of dynamical friction from this point of view. The evolution of the (effective) radius of the perturbation $R$, including angular momentum $L$ and dynamical friction $\eta$, in a $\Lambda$CDM model is given by [64, 65, 66, 67, 68, 69] [^5] In these references, only single parts of the full expression are investigated or reported.